# Fearless Stochasticity in Expectation Propagation

### v1

@@ Appendix C Natural gradient view of EP
$\mu_{i}-\alpha^{-1}\big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\big{]}^{-1}\partialderivative{\lambda_{i}}{\mu_{i}}$ $=\mu_{i}-\alpha^{-1}\big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\big{]}^{-1}\partialderivative{\lambda_{i}}{\mu_{i}}\partialderivative{L}{\lambda_{i}}$
@@
$=\mu_{i}-\bigg{(}\nabla A\Big{(}\theta_{0}+\sum_{j}\lambda_{j}\Big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\bigg{)}$
$=\nabla A\Big{(}\theta_{0}+\sum_{j}\lambda_{j}\Big{)}-\bigg{(}\nabla A\Big{(}\theta_{0}+\sum_{j}\lambda_{j}\Big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\bigg{)}$
@@
where the penultimate equality follows from the definition $\mu_{i}=\mathbb{E}_{\tilde{p}_{i}^{(t)}(z)}[s(z)]$, and the fact that $\lambda_{j}^{(t)}=\lambda_{j}\ \forall\ j$ before the update. All that remains is to show that the right hand side of (7) is equal to $\mathbb{E}_{p_{i}(z)}[s(z)]$;
$\big{(}\nabla A\circ\eta_{i}^{(t)}\big{)}\bigg{(}\lambda_{i}-\alpha\Big{(}\theta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)})\Big{)}\bigg{)}$
@@
Appendix D EP-$\bm{\eta}$ update direction derivation
Using the relation $\eta_{i}=\eta_{i}^{(t)}(\lambda_{i})$, and therefore $\lambda_{i}=\big{(}\eta_{i}^{(t)}\big{)}^{-1}(\eta_{i})$, we have
@@ Appendix D EP-$\bm{\eta}$ update direction derivation
$=-\alpha\Big{(}\partialderivative{\mu_{i}}{\eta_{i}}\tilde{F}_{i}(\mu_{i})\partialderivative{\mu_{i}}{\eta_{i}}^{\top}\Big{)}^{-1}\partialderivative{L}{\lambda_{i}}$
$=-\alpha\Big{(}\partialderivative{\eta_{i}}{\mu_{i}}^{\top}\tilde{F}_{i}(\mu_{i})^{-1}\partialderivative{\eta_{i}}{\mu_{i}}\Big{)}\partialderivative{L}{\lambda_{i}}$
$=-\alpha\Big{(}\partialderivative{\eta_{i}}{\mu_{i}}^{\top}\partialderivative{\mu_{i}}{\eta_{i}}\partialderivative{\eta_{i}}{\mu_{i}}\Big{)}\partialderivative{L}{\lambda_{i}}$
$=-\alpha\partialderivative{\eta_{i}}{\mu_{i}}^{\top}\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (30)
Appendix E EP-$\bm{\mu}$ update derivation
@@
$=\mu_{i}-\epsilon\Big{(}\partialderivative{\eta_{i}}{\mu_{i}}\Big{)}^{-1}\partialderivative{\eta_{i}}{\mu_{i}}\partialderivative{\lambda_{i}}{\eta_{i}}\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}$
$=\mu_{i}-\epsilon\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (31)
Appendix F Bias of EP and EP-$\bm{\mu}$
Let the bias in dimension $d$ of $\mu_{i}^{(t+1)}$, after an update at time $t$, be defined as $E[\mu_{i}^{(t+1)}-\bar{\mu}_{i}^{(t+1)}]_{d}$, where $\bar{\mu}_{i}^{(t+1)}$ is the value of $\mu_{i}^{(t+1)}$ after a noise-free update, and expectation is taken over the sampling distributions of all parallel updates at time $t$. Propositions 2 and 3, which we restate and prove below, summarise the effect of decreasing $\alpha$ on EP and EP-$\mu$.
@@ Appendix F Bias of EP and EP-$\bm{\mu}$
Let $\eta_{i}^{(t)}$ and $\mu_{i}^{(t)}$ be the pre-update natural and mean parameters of $\tilde{p}_{i}{(t)}$ respectively. Furthermore, let
$\mu_{i}^{(t)^{\prime}}$ $=\mu_{i}^{(t)}-\alpha^{-1}\big{[}\tilde{F}_{i}^{(t)}\big{]}^{-1}(\mu_{i})\partialderivative{\lambda_{i}}{\mu_{i}}\partialderivative{L}{\lambda_{i}}+\xi_{i}$
@@
be the mean parameters of $\tilde{p}_{i}^{(t)}$ after an EP inner update (4) but before being mapped to mean parameters of $\tilde{p}_{i}^{(t+1)}$ through map (9). $\xi_{i}$ is some zero-mean noise, e.g. due to MC variation. The second equality follows from the results of Appendix C. Similarly, let $\eta_{i}^{(t)^{\prime}}=\nabla A^{*}(\mu_{i}^{(t)^{\prime}})$ be the corresponding natural parameters. Now apply map (9) to convert to mean parameters of $\tilde{p}_{i}^{(t+1)}$,
$\mu_{i}^{(t+1)}$ $=\Big{(}\nabla A\circ\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\circ\nabla A^{*}\Big{)}\big{(}\mu_{i}^{(t)^{\prime}}\big{)}$
$=\nabla A\bigg{(}\Big{(}\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\Big{)}\Big{(}\nabla A^{*}\big{(}\mu_{i}^{(t)^{\prime}}\big{)}\Big{)}\bigg{)}$
@@
where the last equality follows from the definition of $\eta_{i}^{(t)}(.)$, and by observing that
@@
for all $j$. Let $\bar{\mu}_{i}^{(t)^{\prime}}=\mathbb{E}_{p_{i}(z)}[s(z)]$ be the mean parameters of $\tilde{p}_{i}^{(t)}$ after a noise-free update (but before mapping to those of $\tilde{p}_{i}^{(t+1)}$), so that $\mu_{i}^{(t)^{\prime}}=\bar{\mu}_{i}^{(t)^{\prime}}+\xi_{i}$. Substituting this in, we have
@@
Now subtract the noise-free update $\bar{\mu}_{i}^{(t+1)}$, and take expectations with respect to $\xi_{j}$ for $j=1,\ldots,m$,
@@
$-\alpha\Big{(}\theta_{0}+\sum_{j\neq i}\lambda_{j}-\nabla A\big{(}E_{p_{i}(z)}[s(z)]+\xi_{i}\big{)}\Big{)}^{\top}\partialderivative{L}{\lambda_{i}}+O(\alpha^{2}),$ (38)
which is clearly $O(\alpha)$ as $\alpha\rightarrow^{+}$. ∎
@@
Let $\eta_{i}^{(t)}$ and $\mu_{i}^{(t)}$ be the pre-update natural and mean parameters of $\tilde{p}_{i}{(t)}$ respectively. Furthermore, let
$\mu_{i}^{(t)^{\prime}}$ $=\mu_{i}^{(t)}-\epsilon\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]-\xi_{i}\Big{)}$
@@
be the mean parameters of $\tilde{p}_{i}^{(t)}$ after an EP-$\mu$ update (12) but before being mapped to mean parameters of $\tilde{p}_{i}^{(t)}$ through (9). $\xi_{i}$ is some zero-mean noise, e.g. due to MC variation. The second equality follows from the definition $\mu_{i}^{(t)}=\mathbb{E}_{\tilde{p}_{i}^{(t)}(z)}[s(z)]$, and the fact that $\tilde{p}_{i}^{(t)}=p$ before an update. Now apply map (9) to convert to mean parameters of $\tilde{p}_{i}^{(t+1)}$,
$\mu_{i}^{(t+1)}$ $=\Big{(}\nabla A\circ\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\circ\nabla A^{*}\Big{)}\big{(}\mu_{i}^{(t)^{\prime}}\big{)}$
$=\nabla A\bigg{(}\Big{(}\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\Big{)}\Big{(}\nabla A^{*}\big{(}\mu_{i}^{(t)^{\prime}}\big{)}\Big{)}\bigg{)}$
$=\nabla A\bigg{(}\nabla A^{*}\big{(}\mu_{i}^{(t)}\big{)}+\sum_{j}\Big{(}\nabla A^{*}\big{(}\mu_{j}^{(t)^{\prime}}\big{)}-\nabla A^{*}\big{(}\mu_{j}^{(t)}\big{)}\Big{)}\bigg{)},$ (40)
where the last equality is obtained from the definition of $\eta_{i}^{(t)}(.)$ for $\alpha=1$, and using
@@
$=A^{*}\big{(}\mu_{j}^{(t)^{\prime}}\big{)}-\nabla A^{*}\big{(}\mu_{j}^{(t)}\big{)},$ (41)
for all $j$. Substituting in the definition of $\mu_{j}^{(t)^{\prime}}$, we have
@@
Taking Taylor expansions around $\epsilon=0$, we have
@@
Finally, subtract the noise free $\bar{\mu}_{i}^{(t+1)}$, and take expectations (over $\xi_{j}$ for $j=1,\ldots,m$) to obtain the bias
@@
Appendix G Computational cost analysis

### v2

@@ Appendix D Natural gradient view of EP
$\mu_{i}-\alpha^{-1}\big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\big{]}^{-1}\partialderivative{L_{i}}{\mu_{i}}$ $=\mu_{i}-\alpha^{-1}\big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\big{]}^{-1}\partialderivative{\lambda_{i}}{\mu_{i}}\partialderivative{L}{\lambda_{i}}$
@@
$=\mu_{i}-\left(\nabla A\Big{(}\eta_{0}+\sum_{j}\lambda_{j}\Big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\right)$
$=\mu_{i}-\left(\mu_{i}-\mathbb{E}_{p_{i}(z)}[s(z)]\right)$
@@
where the penultimate equality follows from the definition $\smash{\mu_{i}=\mathbb{E}_{\cramped{\tilde{p}_{i}}^{(t)}(z)}[s(z)]}$, and the fact that $\smash{\cramped{\lambda_{j}}^{(t)}=\lambda_{j}\ \forall\ j}$ before the update. All that remains is to show that the right hand side of (7) is equal to $\mathbb{E}_{p_{i}(z)}[s(z)]$.
$\big{(}\nabla A\circ\eta_{i}^{(t)}\big{)}\bigg{(}\lambda_{i}-\alpha\Big{(}\eta_{0}+\sum_{j}\lambda_{j}-\nabla A^{*}\bm{(}\mathbb{E}_{p_{i}(z)}[s(z)]\bm{)})\Big{)}\bigg{)}$
@@
Appendix E EP-$\bm{\eta}$ update direction derivation
Using the relation $\smash{\eta_{i}=\cramped{\eta_{i}}^{(t)}(\lambda_{i})}$, and therefore $\smash{\lambda_{i}=\cramped{(\cramped{\eta_{i}}^{(t)})}^{-1}(\eta_{i})}$, we have
@@ Appendix E EP-$\bm{\eta}$ update direction derivation
$=-\alpha\Big{(}\partialderivative{\mu_{i}}{\eta_{i}}\tilde{F}_{i}(\mu_{i})\partialderivative{\mu_{i}}{\eta_{i}}\Big{)}^{-1}\partialderivative{L}{\lambda_{i}}$
$=-\alpha\Big{(}\partialderivative{\eta_{i}}{\mu_{i}}\tilde{F}_{i}(\mu_{i})^{-1}\partialderivative{\eta_{i}}{\mu_{i}}\Big{)}\partialderivative{L}{\lambda_{i}}$
$=-\alpha\Big{(}\partialderivative{\eta_{i}}{\mu_{i}}\partialderivative{\mu_{i}}{\eta_{i}}\partialderivative{\eta_{i}}{\mu_{i}}\Big{)}\partialderivative{L}{\lambda_{i}}$
$=-\alpha\partialderivative{\eta_{i}}{\mu_{i}}\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (41)
Appendix F EP-$\bm{\mu}$ update derivation
@@
$=\mu_{i}-\epsilon\Big{(}\partialderivative{\eta_{i}}{\mu_{i}}\Big{)}^{-1}\partialderivative{\eta_{i}}{\mu_{i}}\partialderivative{\lambda_{i}}{\eta_{i}}\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}$
$=\mu_{i}-\epsilon\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (42)
Appendix G Bias of EP and EP-$\bm{\mu}$
Let the bias in dimension $d$ of $\smash{\cramped{\mu_{i}}^{(t+1)}}$, after an update at time $t$, be defined as $\smash{\mathbb{E}[\cramped{\mu_{i}}^{(t+1)}-\cramped{\bar{\mu}_{i}}^{(t+1)}]_{d}}$, where $\smash{\cramped{\bar{\mu}_{i}}^{(t+1)}}$ is the value of $\smash{\cramped{\mu_{i}}^{(t+1)}}$ after a noise-free update, and expectation is taken over the sampling distributions of all parallel updates at time $t$. Propositions 2 and 3, which we restate and prove below, summarise the effect of decreasing $\alpha$ and $\epsilon$ on EP and EP-$\mu$, respectively.
@@ Appendix G Bias of EP and EP-$\bm{\mu}$
Let $\smash{\cramped{\eta_{i}}^{(t)}}$ and $\smash{\cramped{\mu_{i}}^{(t)}}$ be the pre-update natural and mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ respectively. Furthermore, let
$\mu_{i}^{\prime(t)}$ $=\mu_{i}^{(t)}-\alpha^{-1}\big{[}\tilde{F}_{i}^{(t)}(\mu_{i})\big{]}^{-1}\partialderivative{\lambda_{i}}{\mu_{i}}\partialderivative{L}{\lambda_{i}}+\xi_{i}$
@@
be the mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ after an EP inner update (4) but before being mapped to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t+1)}}$ through map (9). $\xi_{i}$ is some zero-mean noise, e.g. due to MC variation. The second equality follows from the results of Appendix D. Similarly, let $\smash{\cramped{\eta_{i}^{\prime}}^{(t)}=\nabla A^{*}(\cramped{\mu_{i}^{\prime}}^{(t)})}$ be the corresponding natural parameters. Now apply map (9) to convert to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t+1)}}$,
$\mu_{i}^{(t+1)}$ $=\Big{(}\nabla A\circ\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\circ\nabla A^{*}\Big{)}\big{(}\mu_{i}^{\prime(t)}\big{)}$
$=\nabla A\bigg{(}\Big{(}\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\Big{)}\Big{(}\nabla A^{*}\big{(}\mu_{i}^{\prime(t)}\big{)}\Big{)}\bigg{)}$
@@
where the last equality follows from the definition of $\smash{\cramped{\eta_{i}}^{(t)}(.)}$, and by observing that
@@
for all $j$. Let $\smash{\cramped{\bar{\mu}_{i}^{\prime}}^{(t)}}$ be the mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ after a noise-free update, but before mapping to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t+1)}}$, so that $\smash{\cramped{\mu_{i}^{\prime}}^{(t)}=\cramped{\bar{\mu}_{i}^{\prime}}^{(t)}+\xi_{i}}$. Substituting this in, we have
@@
Now subtract the noise-free update $\smash{\cramped{\bar{\mu}_{i}}^{(t+1)}}$, and take expectations with respect to $\xi_{j}$ for $j=1,\ldots,m$,
@@
$-\alpha\Big{(}\eta_{0}+\sum_{j\neq i}\lambda_{j}-\nabla A\big{(}E_{p_{i}(z)}[s(z)]+\xi_{i}\big{)}\Big{)}^{\top}\partialderivative{L}{\lambda_{i}}+O(\alpha^{2}),$ (49)
which is clearly $O(\alpha)$ as $\alpha\rightarrow 0^{+}$. ∎
@@
Let $\smash{\cramped{\eta_{i}}^{(t)}}$ and $\smash{\cramped{\mu_{i}}^{(t)}}$ be the pre-update natural and mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ respectively. Furthermore, let
$\mu_{i}^{\prime(t)}$ $=\mu_{i}^{(t)}-\epsilon\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]-\xi_{i}\Big{)}$
@@
be the mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ after an EP-$\mu$ update (12) but before being mapped to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ through (9). $\xi_{i}$ is some zero-mean noise, e.g. due to MC variation. The second equality follows from the definition $\smash{\cramped{\mu_{i}}^{(t)}=\mathbb{E}_{\cramped{\tilde{p}_{i}}^{(t)}(z)}[s(z)]}$, and the fact that $\smash{\cramped{\tilde{p}_{i}}^{(t)}=p}$ before an update. Now apply map (9) to convert to mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t+1)}}$,
$\mu_{i}^{(t+1)}$ $=\Big{(}\nabla A\circ\eta_{i}^{(t+1)}\circ\big{(}\eta_{i}^{(t)}\big{)}^{-1}\circ\nabla A^{*}\Big{)}\left(\mu_{i}^{\prime(t)}\right)$
$=\nabla A\bigg{(}\Big{(}\eta_{i}^{(t+1)}\circ\left(\eta_{i}^{(t)}\right)^{-1}\Big{)}\Big{(}\nabla A^{*}\left(\mu_{i}^{\prime(t)}\right)\Big{)}\bigg{)}$
$=\nabla A\bigg{(}\nabla A^{*}\left(\mu_{i}^{(t)}\right)+\sum_{j}\Big{(}\nabla A^{*}\left(\mu_{j}^{(t)^{\prime}}\right)-\nabla A^{*}\left(\mu_{j}^{(t)}\right)\Big{)}\bigg{)},$ (51)
where the last equality is obtained from the definition of $\smash{\cramped{\eta_{i}}^{(t)}(.)}$ for $\alpha=1$, and using
@@
$=A^{*}\big{(}\mu_{j}^{(t)^{\prime}}\big{)}-\nabla A^{*}\big{(}\mu_{j}^{(t)}\big{)}$ (52)
for all $j$. Substituting in the definition of $\smash{\cramped{\mu_{j}^{\prime}}^{(t)}}$, we have
@@
Taking a Taylor expansion around $\epsilon=0$, we have
@@
Finally, subtract the noise free $\smash{\cramped{\bar{\mu}_{i}}^{(t+1)}}$, and take expectations (over $\xi_{j}$ for $j=1,\ldots,m$) to obtain the bias
@@
Appendix H Implicit geometries of EP variants
In this paper we have shown that the updates of EP, EP-$\eta$, and EP-$\mu$ can be interpreted as performing NGD of $L$ with respect to the distributions $\smash{\{\cramped{p_{i}}^{(t)}\}_{i}}$. The SNEP method of Hasenclever et al. [17], in contrast, performs NGD of $L$ with respect to the site potentials, treating them as bona fide distributions in $\mathcal{F}$ for the purposes of NGD – that is, NGD is performed with respect to the distributions with densities given by $\exp\bm{(}\lambda_{i}^{\top}s(z)-A(\lambda_{i})\bm{)}$ for $i=1,\ldots,m$. In Section 3.4, we argued that the statistical manifolds of $\smash{\{\cramped{p_{i}}^{(t)}\}_{i}}$ are far more relevant for the optimisation of $L$ than those of the site potential pseudo-distributions. In this appendix we shall justify this claim.
The loss function $L$, given by (3), has a unique optimum with respect to $\{\lambda_{i}\}_{i}$ when $\nabla_{\lambda_{i}}L(\theta,\lambda_{1},\ldots,\lambda_{m})=0\;\forall\;i$, which implies that
$\nabla A(\eta_{0}+\sum_{j}\lambda_{j})$ $=-\beta_{i}\nabla_{\lambda_{i}}A_{i}(\theta-\beta^{-1}\lambda_{i},\beta^{-1})\Leftrightarrow$
$\mathbb{E}_{p(x)}[s(z)]$ $=\mathbb{E}_{p_{i}(x)}[s(z)].$ (56)
The minimisation of $L$ with respect to $\{\lambda_{i}\}_{i}$ can therefore be viewed as a joint optimisation in the $m+1$ distributions $p,p_{1},\;\ldots,\;p_{m}$, with the minimum found when their expected $\mathcal{F}$-statistics match. Intuitively, it seems that optimisation of $L$ with respect to $\lambda_{i}$ should be performed with consideration for the geometry of $p$ and $p_{i}$. By instead performing NGD with respect to the distributions $\smash{\{\cramped{p_{i}}^{(t)}\}_{i}}$, in the case of EP, EP-$\eta$ and EP-$\mu$, or with respect to the site potential psuedo-distributions, in the case of SNEP, these methods can be interpreted as using surrogate distributions for NGD [46]. It is possible to show that the Fisher information matrix (FIM) of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ with respect to $\lambda_{i}$ is equal to that of $p$, up to a scalar constant, and furthermore, it can also be seen as a close approximation to that of $p_{i}$. This motivates the use of NGD with respect to $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ for the optimisation of $L$. In contrast, the FIMs of the site-potential pseudo-distributions, used by SNEP, can bear little relation to those of $p$ or $p_{i}$. In particular, the difference will be greatest when the full approximation $\smash{\eta_{0}+\cramped{\sum}_{j}\lambda_{j}}$ is significantly different than $\lambda_{i}$, and we believe this may be why Hasenclever et al. [17] found that increasing the number of sites reduced the convergence speed of SNEP.
We can make these arguments more concrete by considering the curvature of $L$. Taking second derivatives of $L$ with respect to $\lambda_{i}$, we have
$\nabla_{\lambda_{i}}^{2}L(\theta,\lambda_{1},\ldots,\lambda_{m})$ $=\nabla_{\lambda_{i}}^{2}A(\eta_{0}+\sum_{j}\lambda_{j})+\beta_{i}\nabla_{\lambda_{i}}^{2}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}.$ (57)
Let $\lambda_{j}^{(t)}=\lambda_{j}$ at time $t$ for $j=1,\ldots m$, then
$\nabla_{\lambda_{i}=\lambda_{i}^{(t)}}^{2}L(\theta,\lambda_{1},\ldots,\lambda_{m})$ $=\nabla_{\lambda_{i}}^{2}A(\eta_{0}+\sum_{j}\lambda_{j}^{(t)}+(\lambda_{i}-\lambda_{i}^{(t)}))+\beta_{i}\nabla_{\lambda_{i}}^{2}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}$
$=\partialderivative{\mu_{i}}{\eta_{i}}+\beta_{i}\nabla_{\lambda_{i}}^{2}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)},$ (58)
where $\eta_{i}$ and $\mu_{i}$ are the natural and mean parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$, respectively. The Hessian of the second term will not generally be available in closed form, or easily invertible. We can however find a convenient approximation. Note that after an outer update, at time $t$, we have $\smash{\theta=\eta_{0}+\cramped{\sum}_{j}\cramped{\lambda_{j}}^{(t)}}$. Also we have that $c\exp(\lambda_{i}^{\top}s(z))\approx\exp\bm{(}\ell_{i}(z)\bm{)}$, at least near convergence, where $c$ is some unknown scalar constant. Combining these, we have
$\nabla_{\lambda_{i}}^{2}A_{i}\bm{(}(\theta-\beta_{i}^{-1}\lambda_{i},\beta_{i}^{-1})\bm{)}$ $=\nabla_{\lambda_{i}}^{2}\left\{\log\int\exp\left((\theta-\beta^{-1}\lambda_{i})^{\top}s(z)+\beta^{-1}\ell_{i}(z)\right)\mathrm{d}\nu(z)\right\}$
$\approx\nabla_{\lambda_{i}}^{2}\Big{\{}\log\int$dν(z) }
$=\nabla_{\lambda_{i}}^{2}A\left(\eta_{0}+\sum_{j}\lambda_{j}^{(t)}+\beta^{-1}(\lambda_{i}^{(t)}-\lambda_{i})\right)$
$=\nabla_{\lambda_{i}}^{2}A\left(\eta_{0}+\sum_{j}\lambda_{j}^{(t)}+\beta^{-1}(\lambda_{i}-\lambda_{i}^{(t)})\right)$
$=\beta^{-2}\partialderivative{\mu_{i}}{\eta_{i}}.$ (59)
Combining this with (58), we have
$\nabla_{\lambda_{i}=\lambda_{i}^{(t)}}^{2}L(\theta,\lambda_{1},\ldots,\lambda_{m})$ $\approx(1+\beta_{i}^{-1})\partialderivative{\mu_{i}}{\eta_{i}},$ (60)
and if we use this curvature approximation to perform a quasi-Newton step in $L$ with respect to $\lambda_{i}$ and step size $\tilde{\epsilon}$, we have
$\lambda_{i}$ $\leftarrow\lambda_{i}-\tilde{\epsilon}(1+\beta_{i}^{-1})^{-1}\partialderivative{\eta_{i}}{\mu_{i}}\nabla_{\lambda_{i}}L(\theta,\lambda_{1},\ldots,\lambda_{m})$
$=\lambda_{i}-\tilde{\epsilon}(1+\beta_{i}^{-1})^{-1}\partialderivative{\eta_{i}}{\mu_{i}}\left(\nabla A(\eta_{0}+\sum_{j}\lambda_{j})-\mathbb{E}_{p_{i}(z)}[s(z)]\right).$ (61)
By letting $\epsilon=\tilde{\epsilon}(1+\beta_{i}^{-1})^{-1}$ we have equivalence with the update of EP-$\eta$, given by (11). We therefore have that EP-$\eta$ is equivalent to performing a quasi-Newton step in $L$. Furthermore, the approximate curvature becomes exact at the optimum if $\ell_{i}(z)$ is an affine function of $s(z)$, e.g. if $\exp\bm{(}\ell_{i}(z)\bm{)}$ is an (unnormalised) member of $\mathcal{F}$.
In contrast, let us consider the implicit curvature approximation used by SNEP. Let $\gamma_{i}$ be mean parameters of the member of $\mathcal{F}$ with natural parameter $\lambda_{i}$. To perform the inner minimisation of the variational objective, SNEP performs NGD in $L$ with respect to $\gamma_{i}$. The implicit curvature matrix used by the SNEP update is then the Fisher with respect to $\gamma_{i}$, which is given by $\smash{\partialderivative{\lambda_{i}}{\gamma_{i}}=\nabla A^{*}(\gamma_{i})}$. Now consider the actual curvature of $L$ with respect to $\gamma_{i}$,
$\nabla_{\gamma_{i}}^{2}L(\gamma_{i})$ $=\partialderivative{\lambda_{i}}{\gamma_{i}}\nabla_{\lambda_{i}}^{2}L(\theta,\lambda_{1},\ldots,\lambda_{m})\partialderivative{\lambda_{i}}{\gamma_{i}}+\sum_{k=1}^{d}\left[\partialderivative{L}{\lambda_{i}}\right]_{k}\nabla_{\gamma_{i}}^{2}\lambda_{i}(\gamma_{i}),$ (62)
where we have used $\lambda_{i}(\gamma_{i})=\nabla A^{*}(\gamma_{i})$ to denote the backwards map. Near the optimum we have $\smash{\partialderivative{L}{\lambda_{i}}\approx\mathbf{0}}$, and so the curvature will be approximately equal to the first term as we approach the optimum. We have already shown that
$\nabla_{\lambda_{i}}^{2}L(\theta,\lambda_{1},\ldots,\lambda_{m})$ $\approx(1-\beta_{i}^{-1})\partialderivative{\mu_{i}}{\eta_{i}}=(1-\beta_{i}^{-1})\nabla A(\eta_{0}+\sum_{j}\lambda_{j}),$ (63)
and so the curvature with respect to $\gamma_{i}$ is
$\nabla_{\gamma_{i}}^{2}L(\gamma_{i})$ $\approx(1-\beta_{i}^{-1})\partialderivative{\lambda_{i}}{\gamma_{i}}\nabla^{2}A(\eta_{0}+\sum_{j}\lambda_{j})\partialderivative{\lambda_{i}}{\gamma_{i}}$
$=(1-\beta_{i}^{-1})\partialderivative{\lambda_{i}}{\gamma_{i}}\nabla^{2}A(\eta_{0}+\sum_{j}\lambda_{j})\left(\nabla^{2}A(\lambda_{i})\right)^{-1}.$ (64)
Compare this with the implicit curvature used by SNEP of $\smash{\partialderivative{\lambda_{i}}{\gamma_{i}}}$, and it is clear that the two are proportional only when $\smash{\eta_{0}+\cramped{\sum}_{j}\lambda_{j}=\lambda_{i}}$. This suggests that if there are many sites ($m$ is large), or if the prior is very informative, the curvature matrix implicity used by SNEP is likely to be significantly different from the true curvature, even near the optimum.
Appendix I Computational cost analysis